# AdaptNC: Adaptive Nonconformity Scores for Uncertainty-Aware Autonomous Systems in Dynamic Environments

### v1

@@ D.2.1 Environment Dynamics
D.2.2 Access Point Geometry
@@
D.2.3 RSSI Observation Model
@@
The path loss component follows the standard log-distance path loss model (Rappaport, 2010). Here, $P_{0}=-30$ dB denotes the reference received power at unit distance, $n=2.2$ is the path loss exponent, and
@@
Large-scale shadowing is modeled as a temporally correlated log-normal process (Gudmundson, 1991). For each access point $i$, the shadowing term evolves according to a first-order Gauss–Markov model:
@@
Small-scale fading is modeled using a Rayleigh fading process generated via a sum-of-sinusoids (Jakes & Cox, 1994) method. For each access point, a complex channel coefficient $h_{t}^{(i)}$ is generated, and the fading contribution is computed as
@@
where $f_{c}=2.4$ GHz is the carrier frequency and $c$ is the speed of light. Fading processes are independent across access points. The Rayleigh distribution of $|h_{t}^{(i)}|$ follows from the assumption of rich multipath with no dominant line-of-sight component (Clarke, 1968; Rappaport, 2010).
D.2.4 Trajectory Predictor
@@ D.4 Multirotor Tracking
D.4.1 Environment with Actuator Degradation
The multirotor state at time $t$ is
$\mathbf{s}_{t}=[x_{t},\dot{x}_{t},y_{t},\dot{y}_{t},z_{t},\dot{z}_{t},\psi_{t},\dot{\psi}_{t},\theta_{t},\dot{\theta}_{t},\delta_{t},\dot{\delta}_{t}]^{\top}\in\mathbb{R}^{12},$
where $(x,y,z)$ denote position, $(\dot{x},\dot{y},\dot{z})$ translational velocity, $\psi$ yaw, $\theta$ pitch, and $\delta$ roll.
The system evolves in discrete time with integration step $\Delta t=0.1$ s using forward Euler integration.
@@ D.4.1 Environment with Actuator Degradation
The control input is
$\mathbf{u}_{t}=[u_{1,t},u_{2,t},u_{3,t},u_{4,t}]^{\top},$
where $u_{1}$ corresponds to collective thrust, $u_{2}$ to pitch torque, $u_{3}$ to roll torque, and $u_{4}$ to yaw torque.
@@
Each actuator is subject to stochastic degradation modeled by a health variable
$\boldsymbol{\Gamma}_{t}\in[0,1]^{4},$
where $\Gamma_{t}^{(i)}=1$ denotes a healthy actuator. The effective control applied to the system is
$\mathbf{u}_{t}^{\text{eff}}=\boldsymbol{\Gamma}_{t}\odot\mathbf{u}_{t}.$
Actuator health evolves according to an additive Wiener process with negative drift:
@@
Translational accelerations are given by
$\ddot{x}_{t}=g\theta_{t},\quad\ddot{y}_{t}=-g\delta_{t},\quad\ddot{z}_{t}=u_{1,t}^{\text{eff}}-g,$
with gravitational constant $g=9.81$ m/s2. Rotational accelerations satisfy
$\ddot{\psi}_{t}=u_{4,t}^{\text{eff}},\quad\ddot{\theta}_{t}=u_{2,t}^{\text{eff}},\quad\ddot{\delta}_{t}=u_{3,t}^{\text{eff}}.$
Yaw is wrapped to $\pi,\pi]$, and pitch and roll are clipped to [0.3,0.3][-0.3,0.3] radians.
D.4.2 Trajectory Predictor

### v2

@@ E.2.1 Environment Dynamics
E.2.2 Access Point Geometry
@@
E.2.3 RSSI Observation Model
@@
The path loss component follows the standard log-distance path loss model [22]. Here, $P_{0}=-30$ dB denotes the reference received power at unit distance, $n=2.2$ is the path loss exponent, and
@@
Large-scale shadowing is modeled as a temporally correlated log-normal process [13]. For each access point $i$, the shadowing term evolves according to a first-order Gauss–Markov model:
@@
Small-scale fading is modeled using a Rayleigh fading process generated via a sum-of-sinusoids [14] method. For each access point, a complex channel coefficient $h_{t}^{(i)}$ is generated, and the fading contribution is computed as
@@
where $f_{c}=2.4$ GHz is the carrier frequency and $c$ is the speed of light. Fading processes are independent across access points. The Rayleigh distribution of $|h_{t}^{(i)}|$ follows from the assumption of rich multipath with no dominant line-of-sight component [8, 22].
E.2.4 Trajectory Predictor
@@ E.4 Multirotor Tracking
E.4.1 Environment with Actuator Degradation
The multirotor state at time $t$ is: $\mathbf{s}_{t}=[x_{t},\dot{x}_{t},y_{t},\dot{y}_{t},z_{t},\dot{z}_{t},\psi_{t},\dot{\psi}_{t},\theta_{t},\dot{\theta}_{t},\delta_{t},\dot{\delta}_{t}]^{\top}\in\mathbb{R}^{12}$, where $(x,y,z)$ denote position, $(\dot{x},\dot{y},\dot{z})$ translational velocity, $\psi$ yaw, $\theta$ pitch, and $\delta$ roll. The system evolves in discrete time with integration step $\Delta t=0.1$ s using forward Euler integration.
@@ E.4.1 Environment with Actuator Degradation
The control input is: $\mathbf{u}_{t}=[u_{1,t},u_{2,t},u_{3,t},u_{4,t}]^{\top}$, where $u_{1}$ corresponds to collective thrust, $u_{2}$ to pitch torque, $u_{3}$ to roll torque, and $u_{4}$ to yaw torque.
@@
Each actuator is subject to stochastic degradation modeled by a health variable, $\boldsymbol{\Gamma}_{t}\in[0,1]^{4},$ where $\Gamma_{t}^{(i)}=1$ denotes a healthy actuator. The effective control applied to the system is: $\mathbf{u}_{t}^{\text{eff}}=\boldsymbol{\Gamma}_{t}\odot\mathbf{u}_{t}.$ Actuator health evolves according to an additive Wiener process with negative drift:
@@
Translational accelerations are given by $\ddot{x}_{t}=g\theta_{t},~\ddot{y}_{t}=-g\delta_{t},~\ddot{z}_{t}=u_{1,t}^{\text{eff}}-g,$ with gravitational constant $g=9.81$ m/s2. Rotational accelerations satisfy: $\ddot{\psi}_{t}=u_{4,t}^{\text{eff}},~\ddot{\theta}_{t}=u_{2,t}^{\text{eff}},~\ddot{\delta}_{t}=u_{3,t}^{\text{eff}}.$ Yaw is wrapped to $\pi,\pi]$, and pitch and roll are clipped to [0.3,0.3][-0.3,0.3] radians.
E.4.2 Trajectory Predictor
@@ E.4.4 Expert Weights
E.5 Hyperparameter Sensitivity Analysis
Social Navigation
Varying window size (update period = 100)
Window Size Global Cov. ($\uparrow$) Mean Volume ($\downarrow$)
35 0.9043 $\pm$ 0.0144 40.24 $\pm$ 5.66
75 (base) 0.9223 $\pm$ 0.0090 42.00 $\pm$ 5.54
120 0.9641 $\pm$ 0.0043 52.41 $\pm$ 5.46
Varying update period (window size = 75)
Update Period Global Cov. ($\uparrow$) Mean Volume ($\downarrow$)
70 0.9609 $\pm$ 0.0032 36.45 $\pm$ 3.97
100 (base) 0.9223 $\pm$ 0.0090 42.00 $\pm$ 5.54
130 0.8780 $\pm$ 0.0111 46.74 $\pm$ 5.93
Multirotor Tracking
Varying window size (update period = 100)
Window Size Global Cov. ($\uparrow$) Mean Volume ($\downarrow$)
35 0.9180 $\pm$ 0.0050 0.04 $\pm$ 0.01
65 (base) 0.9264 $\pm$ 0.0055 0.04 $\pm$ 0.01
100 0.9620 $\pm$ 0.0049 0.03 $\pm$ 0.01
Varying update period (window size = 65)
Update Period Global Cov. ($\uparrow$) Mean Volume ($\downarrow$)
60 0.9685 $\pm$ 0.0027 0.03 $\pm$ 0.01
100 (base) 0.9264 $\pm$ 0.0055 0.04 $\pm$ 0.01
140 0.8976 $\pm$ 0.0060 0.04 $\pm$ 0.01
Indoor Localization
Varying window size (update period = 20)
Window Size Global Cov. ($\uparrow$) Mean Volume ($\downarrow$)
60 0.9135 $\pm$ 0.0027 26.29 $\pm$ 1.64
100 (base) 0.9062 $\pm$ 0.0027 27.81 $\pm$ 1.75
140 0.9004 $\pm$ 0.0032 28.67 $\pm$ 1.71
Varying update period (window size = 100)
Update Period Global Cov. ($\uparrow$) Mean Volume ($\downarrow$)
15 0.8964 $\pm$ 0.0034 25.00 $\pm$ 1.71
20 (base) 0.9062 $\pm$ 0.0027 27.81 $\pm$ 1.75
50 0.9365 $\pm$ 0.0024 35.90 $\pm$ 2.61
Table 5: Sensitivity analysis with respect to window size and update period across all environments. Performance remains stable around the chosen base configurations, indicating robustness to moderate hyperparameter variations.